OpenStudy (anonymous):

For a given geometric sequence, the 5th term, a[5] , is equal to 11/16 , and the 8th term, a[8] , is equal to 44 . Find the value of the 11th term, a[11] . If applicable, write your answer as a fraction.

OpenStudy (anonymous):

Please help quickly!!!

OpenStudy (anonymous):

this is not exactly on point but its close. 5459/50 or 109.18.

OpenStudy (campbell_st):

well a term in a geometric sequence is \[a_{n} = a_{1} \times r^{n -1}\] so find the common difference \[\frac{11}{16} = a_{1} \times r^{5-1}\] and \[44 = a_{1} \times r^{8-1}\] rewriting the equation gives \[44 = a_{1} \times r^4 \times r^3\] now subtituting the 5th term \[44 = \frac{11}{16} \times r^3\] then \[64 = r^3\] so the common ratio is r = 4 since you know the 8th term the 11th term will be \[a_{11} = a_{1} \times r^{11 -1}....or.... a_{1} \times r^{10}\] rewriting in terms of the 8th term \[a_{11} = a_{1} \times r^3 \times r^3\] so \[a_{11} = 44 \times 4^3\] or \[a_{11} = 44 \times 64\]

OpenStudy (anonymous):

that ain't right.

OpenStudy (anonymous):

oh wait. I was adding.

OpenStudy (campbell_st):

ok well have a look at this 5th term 11/16 6th term 11/16 x 4 = 11/4 7th term 11/4 x 4 = 11 8th term 11 x 4 = 44 funny I seem to have the common ratio correct... 9th term 44 x 4 = 176 10th term 176 x 4 = 704 11th term 704 x 4 = 2816 perhaps you could post your solution...it would make interesting reading...